OpenStudy (anonymous):

Is there a slick/quick/better way to get the \(nth\) derivative of \(x f(x) = \log x\)? Or, to show \[f^{(n)}(1) = (-1)^{n+1} n! \left(1 + \frac12 + \ldots +\frac1n\right)\]

OpenStudy (kinggeorge):

Perhaps induction?

OpenStudy (anonymous):

Hmm how do I use induction on this? I don't even know the pattern. Is differentiating and trying to make out some pattern the only way? Here's the actual question.

OpenStudy (kinggeorge):

First, rearrange to get \(f(x)=\log(x)/x\), and then take the first one or two or three or four derivatives. From this you might be able to find some formula for the \(n\)-th derivative, and from that find \[f^{(n)} (1)\]

OpenStudy (kinggeorge):

Looking at Wolfram, it seems like the \(n\)-th derivative would be \[f^{(n)} = (-1)^n \cdot {n! \log(x) - (2n! -1) \over x^{n+1}}\]And it looks possible to prove this with induction.

OpenStudy (anonymous):

Thanks, I will see what I can get. I did try differentiating earlier but wasn't able to make out the pattern.

OpenStudy (asnaseer):

I don't know if this helps, but you could try implicit differentiation:\[\begin{align} xf(x)&=\log(x)\\ xf^1(x)+f(x)&=\frac{1}{x}\\ xf^2(x)+f^1(x)+f^1(x)&=-\frac{1}{x^2}\\ xf^2(x)+2f^1(x)&=-\frac{1}{x^2}\\ ... \end{align}\] this eventually seems to lead to the following pattern:\[xf^n(x)+nf^{n-1}(x)=(-1)^{n+1)*\frac{n-1}{x^n}\]

OpenStudy (asnaseer):

the rendering of the last line hasn't come out right. it should be:\[xf^n(x)+nf^{n-1}(x)=(-1)^{n+1}*\frac{n-1}{x^n}\]

OpenStudy (asnaseer):

so setting x=1 leads to:\[f^n(1)+nf^{n-1}(1)=(-1)^{n+1}*(n-1)\]

OpenStudy (kinggeorge):

We would still need to solve for get rid of the recurrence relation though. And that may or may not be easier.

OpenStudy (asnaseer):

true...

OpenStudy (kinggeorge):

I didn't type that sentence very well :(

OpenStudy (asnaseer):

it's lucky this isn't the English or Writing group then or you would be real trouble ;-)

OpenStudy (asnaseer):

*be in

OpenStudy (anonymous):

I did try that too but I wasn't able to get it all together. :/

OpenStudy (asnaseer):

maybe you can use the recurrence result to prove KingGeorge's result above using induction?

OpenStudy (asnaseer):

hmmm... but even with that I can't see how it would lead to:\[f^{(n)}(1) = (-1)^{n+1} n! \left(1 + \frac12 + \ldots +\frac1n\right)\]

OpenStudy (anonymous):

I am trying it's only the algebra now, with enough trials I might get it.

OpenStudy (asnaseer):

ok - good luck! :)

OpenStudy (kinggeorge):

I went a few terms farther with my possible definition of \(f^{(n)}\) and it doesn't seem to work anymore.

OpenStudy (asnaseer):

I worked with the recurrence relation and ended up with something close (but no cigar!):\[f^n(1)=(-1)^{n+1}(\frac{1}{n(n-2)(n-3)...}+\frac{1}{(n-1)(n-3)(n-4)...}+...+1)\]

OpenStudy (asnaseer):

sorry that should be:\[f^n(1)=(-1)^{n+1}n!(\frac{1}{n(n-2)(n-3)...}+\frac{1}{(n-1)(n-3)(n-4)...}+...+1)\]

OpenStudy (asnaseer):

I'm guessing I either made a mistake in the algebra somewhere or the derivation of the recurrence relation had an error in it.

OpenStudy (anonymous):

Here's what I am getting \[xf^n(x) = \small (-1)^n\cdot \frac 1{x^n}\cdot \left((n-1) + n(n-2) + n(n-1)(n-3) + n(n-1)(n-2)(n-4) + \ldots \right)\]

OpenStudy (asnaseer):

yes Ishaan94 - if you take an n! out that you will get the same as I did

OpenStudy (asnaseer):

you can set x=1 to simplify the algebra

OpenStudy (kinggeorge):

This may or may not be helpful, but the new possibility I have for the \(n\)-th derivative is \[\large f^{(n)}(x) = (-1)^n \cdot {n! \log(x) - n!({1 \over 1}+{1 \over 2}+{1 \over 3}+...+{1 \over n}) \over x^{n+1}} \]Which would get you what you need if you factor a \(-1\) out of the top. Proving this, is still not easy.

OpenStudy (asnaseer):

does this lead directly from your previous result?

OpenStudy (kinggeorge):

My previous result was flawed. After finding some more patterns however, this seems correct. It works perfectly through the 5th derivative.

OpenStudy (asnaseer):

ok - then this looks a lot more promising - great work!

OpenStudy (asnaseer):

KingGeorge - why do we need to prove this - I thought you said it comes from repeated derivatives and then spotting the pattern? isn't that proof enough?

OpenStudy (anonymous):

I don't understand why didn't we get the answer even after solving the recurrence? Nice work, KingGeorge :-)

OpenStudy (kinggeorge):

We still need to prove the pattern works for all \(n\). And if I'm doing my derivatives correctly, I think this is straightforward to prove using induction.

OpenStudy (kinggeorge):

It uses a lot of paper, but it's straightforward.

OpenStudy (asnaseer):

ok - I guess you now have a way forward now Ishaan94 (thanks to the King!) :)

OpenStudy (anonymous):

Awesome, it's perfect. Thanks asnaseer and KingGeorge :-) #Sorry for replying late, I was afk.

OpenStudy (kinggeorge):

In the derivative, the only part that's a little hard is proving that \[n!+n!\left(1+{1 \over2}+...+{1 \over n}\right) = (n+1)!\left(1+{1 \over 2}+...+{1 \over n+1}\right)\]The rest is rather easy.

OpenStudy (anonymous):

I did prove it and it works perfectly.

OpenStudy (anonymous):

Thanks again. and great work. :D

OpenStudy (kinggeorge):

You're welcome.